A 33cm -long solenoid, 2.0cm in diameter, is to produce a 0.25T magnetic field at its center.

If the maximum current is 4.2A , how many turns must the solenoid have?
Answer #1

B due to solonid is B = u0nI

n = 0.25/(4pi*10^-7 * 4.2)

n = 47391

N = nl = 47391*0.33 = 15640 turns